Royal Minty of Britain has purchased 20,000 ounces of silver from Silver Products at US$8.30, payable in 180 days. The current spot rate is 1.8127 ($US/£) and the 180-day forward is 1.7863. The CEO at Royal Minty suggests that the spot rate in six months time will be 1.7915.

Interest rates in Britain are currently 4.70 percent for 180 days and 1.15 percent in the United States.

a-1. Calculate the receipts if Royal Minty takes a chance on the spot rate. (Round the final answer to the nearest whole pound.)

Receipts           £ 92660 92660 Correct

  

a-2. Calculate the receipts if Royal Minty books a forward contract. (Round the final answer to the nearest whole pound.)

Receipts           £ 92930 92930 Correct

    

a-3. Calculate the receipts if Royal Minty buys a money market hedge. (Round intermediate calculations and the final answer to the nearest whole pound.)

Receipts           £ 94790 94790 Incorrect  
Answer #1

Cost of Silver = 20000*8.30 = $ 166,000

Answer a-1)  the receipts if Royal Minty takes a chance on the spot rate

The exchange rate in spot ($US/£) = 1.7915

Receipts = 166,000 / 1.7915= £ 92659.782 i.e £ 92660

Answer a-2)  the receipts if Royal Minty takes a chance on the forward rate

The exchange rate in forward  ($US/£) = 1.7863

Receipts = 166,000 / 1.7863= £ 92929.519 i.e £ 92930.

Answer a-3)Effective exchange rate in case of Money market Hedge as

Exchange rate = S * (1+i$) / (1+ip ) =1.8127 *1.0115/1.047 = 1.7512

Receipts = 166,000 / 1.7512= £ 94790.0927 i.e £ 94790